OpenStudy (anonymous):

let A = power set of ℝ. Define f: ℝ->A by formula f(x) = {y∈ℝ | y^2 < x}. Is one-to-one? is it onto?

OpenStudy (anonymous):

suppose f(x1) = f(x2). I'm trying to determine whether x1 = x2. So far I got, f(x1) = {y∈ℝ | y^2 < x1} f(x2) = {y∈ℝ | y^2 < x2} the fact that f(x) isn't a real number throws me off, thus, I don't really know how to solve for x1 and x2 and show they are (or are not) equal

OpenStudy (anonymous):

Well, \(f(x)\) is a just a set of real numbers.

OpenStudy (anonymous):

right, f(x) is a set, which isn't a number

OpenStudy (anonymous):

Consider \(x=0\) and consider \(x=-1\). That will give away the answer.

OpenStudy (anonymous):

x = 0 then f(0) is a set of y that y^2 < 0, which is a empty set. x = 1 gives -1 < y < 1

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

I said \(x=-1\) not \(x=1\).

OpenStudy (anonymous):

oh... f(-1) also gives an empty set. Ah so f isn't one-to-one

OpenStudy (anonymous):

The onto part should be pretty obvious as well.

OpenStudy (anonymous):

f is onto right?

OpenStudy (anonymous):

Think about it for a few seconds.

OpenStudy (anonymous):

Consider a function: \[ g(x) = \frac{1}{x-3} \]Would it's domain be in \(A\)? If it is, then \(f(x)\) has to map to it.

OpenStudy (anonymous):

its^

OpenStudy (anonymous):

{3} isn't in A

OpenStudy (anonymous):

First of all, the domain would be \(\mathbb R \setminus \{3\}\).

OpenStudy (anonymous):

Is \(\mathbb R \setminus \{3\}\subseteq \mathbb R\)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

For that matter is \(\{3\}\subseteq \mathbb R\)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait yeah

OpenStudy (anonymous):

How are you defining the power set?

OpenStudy (anonymous):

set of all subsets

OpenStudy (anonymous):

Okay, so why would you think it isn't in \(A\)?

OpenStudy (anonymous):

You said \(A = \mathcal P(\mathbb R)\)

OpenStudy (anonymous):

oh {3} is indeed in A

OpenStudy (anonymous):

Does \(f(x)\) ever map to \(\{3\}\)?

OpenStudy (anonymous):

huhm.. so I need to find an x such that f(x) = {3}, which is impossible. So f is not onto

OpenStudy (anonymous):

I see now. Thank you!

OpenStudy (anonymous):

Can you prove that \(\forall x\quad \{3\}\notin f(x)\)?

OpenStudy (anonymous):

huhm.... suppose x in R, then -sqrt(x) < y < sqrt(x)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

i just solved for y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The way I see it is given, x > 0, f(x) will be a set that has an equal length, and {3} doesn't have a length. So, x > 0 works. Now I need to consider x <= 0, which will gives an empty set

OpenStudy (anonymous):

I see where you are going. Another idea: \(3\in f(x) \implies 3^2<x \implies 2^2<x \implies 2\in f(x)\). So \(\{3\}\notin f(x)\) since \(2\notin \{3\}\).

OpenStudy (anonymous):

you meant to say {3} ∈ ℝ right?

OpenStudy (anonymous):

wait wait typo

OpenStudy (anonymous):

you meant to say {3} ∈ f(x) ?

OpenStudy (anonymous):

I mean to say \(\{3\}\notin \text{Im} [{ f(x)}]\)

OpenStudy (anonymous):

how does {3} ∈ f(x) => 3^2 < x?

OpenStudy (anonymous):

No, it is \(3\in f(x)\).

OpenStudy (anonymous):

It returns a set of real numbers, not a set of sets

OpenStudy (anonymous):

but f(x) is a set...

OpenStudy (anonymous):

Yes, \(f(x)\) is a set. I'm saying suppose that \(3\) is in that set.

OpenStudy (anonymous):

ohh.... :o

OpenStudy (anonymous):

If \(3\) is in the set, then \(3^2<x\). We know that \(2^2<x\) based on this fact. This means \(2\) is also in the set.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

Well, your proof that \(0\) must be in \(f(x)\) or \(f(x)\) must be empty also works.

OpenStudy (anonymous):

you mean the proof I gave that f(x) is either empty or a set that has a length?

OpenStudy (anonymous):

Whoops, I didn't read it thoroughly, I thought you were going to say something like either it is empty or \(0\) is in it.

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

well, I did say that -sqrt(x) < y < sqrt(x), so yeah either f(x) is empty or 0 is in it. I didn't thought of it this way when I said it. All I did was merely solving for y. But now I see the connection to what I did

OpenStudy (anonymous):

once again, thank you for your time ^.^